OpenStudy (anonymous):

Determine value of expression when x=2 y=5 and z=-3 4y-3|x-z^2|

OpenStudy (rulnick):

4(5)-3|2-(-3)^2| = 20-3|2-9| = ...

OpenStudy (anonymous):

Do you add the 3 or subtract the 20?

OpenStudy (rulnick):

First do the 2-9, then take the absolute value, then multiply that by 3. That whole thing will then be subtracted from 20 to get your answer.

OpenStudy (anonymous):

Okay thanks!

OpenStudy (rulnick):

welcome